### v1

# Juan Guillermo Garrido

Abstract
This paper is devoted to study a characterization of (strong) local maximal monotonicity in terms of a property involving the graphical derivative of a set-valued mapping defined on a Hilbert space. As a consequence, a second-order characterization of variational convexity is provided without the assumption of subdifferential continuity.
Keywords:
Variational analysis Local monotonicity Prox regular functionsVariational convexityGeneralized differentiation
MSC:
49J5249J53 47H05
1 Introduction
Local monotonicity often appears in variational analysis. For instance, prox regular functions can be characterized via the local monotonicity of subdifferential mapping (see, e.g., Poliquin1996 ; MR2113863 ). Recently, Rockafellar introduced the concept of variational convexity in finite-dimensional spaces MR3995335 , motivated by the local maximal monotonicity of subdifferential mappings. This concept of functions has gained significant attention in recent years, as it has enabled the generalization of numerical algorithms for solving nonsmooth optimization problems (see, e.g., MR4550949 ; MR3995335 ). In MR4784083 , variational convexity is further studied in general Banach spaces.
Generalized differentiation is a fundamental tool for studying the first-order variational properties of set-valued mappings. In particular, the graphical derivative is used to establish criteria for Lipschitzian properties such as metric regularity/subregularity and strong metric regularity (see, e.g., MR2252233 ; MR3108434 ; MR3705353 ). Furthermore, the graphical derivative can be employed to extend numerical algorithms to nonsmooth settings, as in the case of Newton’s method (see, e.g., MR4252735 ; aragon2024coderivative ). Additionally, the explicit computation of graphical derivatives for certain important cases is an active area of research, as it contributes to the study of the stability of solution mappings (see, e.g., MR3288139 ; MR3485980 ; MR4769808 ). For instance, workable calculus rules for graphical derivatives are provided in MR3428695 ; MR3693740 .
In this paper, we aim to study a characterization of local maximal monotonicity via graphical derivatives for general set-valued mappings in Hilbert spaces, as well as its corresponding counterparts for strong monotonicity and hypomonotonicity. In particular, for strong local maximal monotonicity, we provide several characterizations, all of which depend on the lower-definiteness of the graphical derivative (or strict) but require different additional assumptions, such as local maximal hypomonotonicity or strong metric regularity.
A parallel result using coderivatives can be found in (MR3485980, , Theorem 3.4). Previously, a related result was established in (MR3843842, , Theorem 3.3), where the tilt stability property for prox-regular functions is characterized as the lower definiteness of graphical derivative of the subdifferential mapping. Later, in (gfrerer2025, , Theorem 4.3), second-order characterizations of variational convexity for prox-regular functions were provided, and in the same line, in Rockafellar2025 are also studied these kind of results from a point-based perspective.
The paper is organized as follows: In Section 2, we present the necessary mathematical background and notation for the subsequent sections, along with some preparatory results. Then, in Section 3, we establish the main result of the paper and provide examples to illustrate the necessity of our hypotheses. In Section 4, a characterization of variational convexity is developed through a condition for $f$-attentive graphical derivative of subdifferential mapping, avoiding subdifferential continuity. The remainder of the paper are some concluding remarks.
2 Preliminaries and Auxiliary Results
[…]
Given a set-valued mapping $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$, we define its domain as $\operatorname{dom}F:=\{x\in\mathcal{H}:F(x)\neq\emptyset\}$ and its graph as $\operatorname{gph}F:=\{(x,y)\in\mathcal{H}\times\mathcal{H}:y\in F(x)\}$.
Let us recall that a function $f\colon[0,1]\to\mathcal{H}$ is said to be absolutely continuous if for every $\varepsilon>0$, there is $\delta>0$ such that for all pairwise disjoint set of intervals $(]a_{i},b_{i}[)_{i\in\mathbb{N}}\subset[0,1]$ with $\sum_{j\in\mathbb{N}}b_{j}-a_{j}<\delta$, one has $\sum_{j\in\mathbb{N}}\|f(b_{j})-f(a_{j})\|<\varepsilon$. The following proposition can be found in (MR2168068, , Theorem 2.2.17).
Proposition 1
Consider a function $f\colon[0,1]\to\mathcal{H}$. Suppose that $f$ is absolutely continuous, then $f$ is derivable a.e., $f^{\prime}\in L^{1}([0,1];\mathcal{H})$ and
[…]
Let $\mathcal{K}$ be a metric space. For a given set-valued mapping $F\colon\mathcal{K}\rightrightarrows\mathcal{H}$ we recall the following
$\limsup_{x^{\prime}\to x}F(x^{\prime}):=\{y\in\mathcal{H}:\exists(x_{n},y_{n})\to(x,y)\text{ s.t. }\forall n\in\mathbb{N},(x_{n},y_{n})\in\operatorname{gph}F\},$
and
$\liminf_{x^{\prime}\to x}F(x^{\prime}):=\left\{y\in\mathcal{H}\;\middle|\;\begin{aligned} &\forall(x_{n})\to x,\exists\text{ subsequence }(n_{k})\subset\mathbb{N},\\ &\exists(y_{n_{k}})\to y,\forall k\in\mathbb{N},(x_{n_{k}},y_{n_{k}})\in\operatorname{gph}F\end{aligned}\right\}.$
Some concepts of locally monotonicity are recalled. A set-valued mapping $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ is said to be locally monotone around $(\bar{x},\bar{y})\in\operatorname{gph}F$ if there are neighborhoods $U,V$ of $\bar{x},\bar{y}$ respectively such that
$\langle y-y^{\prime},x-x^{\prime}\rangle\geq 0,\ \forall(x,y),(x^{\prime},y^{\prime})\in\operatorname{gph}F\cap(U\times V).$ (1)
We also say that $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ is strongly locally monotone at $(\bar{x},\bar{y})$ if there is $\kappa>0$ such that $F-\kappa\text{Id}$ is locally monotone at $(\bar{x},\bar{y}-\kappa\bar{x})$. Let us say that $\kappa$ is the modulus of strong local monotonicity. We say $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ is locally hypomonotone at $(\bar{x},\bar{y})$ if there is $\kappa>0$ such that $F+\kappa\text{Id}$ is locally monotone at $(\bar{x},\bar{y}+\kappa\bar{x})$.
The set-valued mapping $F$ is said to be locally maximal monotone at $(\bar{x},\bar{y})$ when $F$ is locally monotone at $(\bar{x},\bar{y})$ and there is a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ such that for every globally monotone operator $S\colon\mathcal{H}\rightrightarrows\mathcal{H}$ such that $\operatorname{gph}F\cap U\times V\subset\operatorname{gph}S$ then $\operatorname{gph}F\cap U\times V=\operatorname{gph}S\cap U\times V$. Similarly, we say that $F$ is locally maximal hypomonotone at $(\bar{x},\bar{y})$ if there is $\rho>0$ such that $F+\rho\text{Id}$ is locally maximal monotone at $(\bar{x},\bar{y}+\rho\bar{x})$.
A different definition of maximality was introduced previously in MR1886225 , which for reflexive Banach spaces it becomes equivalent to the one given here (see, e.g., (MR4769808, , Theorem 8.5)).
We say $F$ is strongly locally maximal monotone at $(\bar{x},\bar{y})$ with modulus $\kappa>0$ if there is a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ such that
$\langle y_{1}-y_{2},x_{1}-x_{2}\rangle\geq\kappa\|x_{1}-x_{2}\|^{2},\forall(x_{1},y_{1}),(x_{2},y_{2})\in(U\times V)\cap\operatorname{gph}F,$
and for every monotone operator $S\colon\mathcal{H}\rightrightarrows\mathcal{H}$ such that $U\times V\cap\operatorname{gph}F\subset\operatorname{gph}S$ we have $U\times V\cap\operatorname{gph}F=U\times V\cap\operatorname{gph}S$.
The following proposition shows that strong local maximal monotonicity can be expressed in terms of local maximal monotonicity.
Proposition 2
Let $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ be a set-valued mapping, $(\bar{x},\bar{y})\in\operatorname{gph}F$ and $\kappa>0$. Then, $F$ is strongly locally maximal monotone at $(\bar{x},\bar{y})$ with modulus $\kappa$ if and only if $F-\kappa\text{Id}$ is locally maximal monotone at $(\bar{x},\bar{y}-\kappa\bar{x})$.
Proof
It follows from (MR4769808, , Proposition 6.7 & Corollary 8.9).
We recall some generalized Lipschitz notions for set-valued mappings.
Definition 1
Let $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ be a set-valued mapping with $(x,y)\in\operatorname{gph}F$. We say
  1. 1)
    $F$ is metrically regular with modulus $\kappa$ at $(\bar{x},\bar{y})$ if there is a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ such that
    $d(x;F^{-1}(y))\leq\kappa\cdot d(y;F(x)),\forall(x,y)\in U\times V.$ (2)
    The infimum of $\kappa$ over all the open sets $U\times V$ satisfying (2) is denoted by $\text{reg}(F;(\bar{x},\bar{y}))$ and is called the modulus of metric regularity.
  2. 2)
    $F$ is strongly metrically subregular at $(x,y)$ if there are $\kappa>0$ and a neighborhood $U$ of $\bar{x}$ such that
    $\|x-\bar{x}\|\leq\kappa\cdot d(\bar{y};F(x)),\forall x\in U.$
  3. 3)
    $F$ is strongly metrically regular at $(\bar{x},\bar{y})$ if there is a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ joint with a Lipschitz function $\theta\colon V\to U$ such that $V\times U\cap\operatorname{gph}F^{-1}=\operatorname{gph}\theta$.
Given a set $C\subset\mathcal{H}$ and $x\in C$, the (Bouligand) tangent cone and the strict tangent cone of $C$ at $x$ are defined (see, e.g., MR2458436 ), respectively, by
$T(x;C):=\limsup_{s\searrow 0}\frac{C-x}{s}\text{ and }\displaystyle T_{\ast}(x;C):=\limsup_{\begin{subarray}{c}y\to x\\ s\searrow 0\end{subarray}}\frac{C-y}{s}.$
For a set-valued mapping $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ and $(x,y)\in\operatorname{gph}F$, for $u\in\mathcal{H}$, we define the graphical derivative of $F$ at $(x,y)$ in the direction $u$ as
$DF(x,y)(u):=\{v\in\mathcal{H}:(u,v)\in T((x,y);\operatorname{gph}F)\},$
and the strict graphical derivative of $F$ at $(x,y)$ in the direction $u$ as
$D_{\ast}F(x,y)(u):=\{v\in\mathcal{H}:(u,v)\in T_{\ast}((x,y);\operatorname{gph}F)\}.$
Definition 2
Let us consider $\rho\in\mathbb{R}$. We say that a set-valued mapping $G\colon\mathcal{H}\rightrightarrows\mathcal{H}$ is $\rho$-lower-definite if for all $(u,v)\in\operatorname{gph}G$, $\langle u,v\rangle\geq\rho\|u\|^{2}$. For a given set-valued mapping $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$, we say $DF$ is $\rho$-lower-definite on $A\subset\mathcal{H}\times\mathcal{H}$ if for all $(x,y)\in A\cap\operatorname{gph}F$, $DF(x,y)$ is $\rho$-lower-definite. When $A$ is the neighborhood of a given point $(\bar{x},\bar{y})\in\operatorname{gph}F$, we say $DF$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$.
For a given real number $\kappa\in\mathbb{R}$, we define the $\kappa$-shifted operator $\mathsf{T}_{\kappa}\colon\mathcal{H}\times\mathcal{H}\to\mathcal{H}\times\mathcal{H}$ given by
$\mathsf{T}_{\kappa}(x,y)=(x,y+\kappa x),$ (3)
which is linear and invertible.
Lemma 1
Consider a set-valued mapping $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ such that $DF$ is $\kappa$-lower-definite for some $\kappa\in\mathbb{R}$ on some set $A\subset\mathcal{H}\times\mathcal{H}$. Then, for all $\gamma\in\mathbb{R}$, $D(F+\gamma\text{Id})$ is $(\kappa+\gamma)$-lower-definite on $\mathsf{T}_{\gamma}(A)$.
Proof
It is not difficult to see that for every $(x,y)\in\mathsf{T}_{\gamma}(A)\cap\operatorname{gph}(F+\gamma\text{Id})$
$D(F+\gamma\text{Id})(x,y)=DF(x,y-\gamma x)+\gamma\text{Id}.$
Then, if $v\in D(F+\gamma\text{Id})(x,y)(u)$ it follows that $v-\gamma u\in DF(x,y-\gamma x)(u)$. Since $(x,y-\gamma x)\in A$, we have $DF(x,y-\gamma x)$ is $\kappa$-lower-definite, then $\langle v-\gamma u,u\rangle\geq\kappa\|u\|^{2}$ and it yields $\langle u,v\rangle\geq(\kappa+\gamma)\|u\|^{2}$, obtaining the desired conclusion.
The following lemma states a chain rule for graphical derivative, between a locally Lipschitz mapping and an absolutely continuous function.
Lemma 2
Let $U\subset\mathcal{H}$ be an open set, $F\colon U\rightrightarrows\mathcal{H}$ be a set-valued mapping, which is locally Lipschitz on its domain and $x\colon[0,1]\to U$ be an absolutely continuous function such that $x([0,1])\subset\operatorname{dom}F$. Then, we have that for almost all $t\in[0,1]$
$DF(x(t))(\dot{x}(t))=\left\{\frac{d}{dt}(F\circ x)(t)\right\}.$
Proof
See (MR2458436, , Theorem 5.3.2).
We recall the following result concerning strong metric subregularity, established under the assumption of the lower-definiteness of the graphical derivative.
Lemma 3
Suppose that $\mathcal{H}$ is finite-dimensional. Let us consider $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ to be a set-valued mapping with $(\bar{x},\bar{y})\in\operatorname{gph}F$ such that $F$ has locally closed graph at $(\bar{x},\bar{y})$ and $DF(\bar{x},\bar{y})$ is $\rho$-lower-definite. Then $F$ is strongly metrically subregular at $(\bar{x},\bar{y})$. In particular, $\bar{x}$ is an isolated point in $F^{-1}(\bar{y})$.
Proof
It is a direct consequence of (MR3288139, , Theorem 4E.1).
The next lemma states the lower-definiteness of the strict graphical derivative derived from the local monotonicity.
Lemma 4
Consider a set-valued mapping $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ and $(\bar{x},\bar{y})\in\operatorname{gph}F$. Suppose that $F-\kappa\text{Id}$ is locally monotone at $(\bar{x},\bar{y}-\kappa\bar{x})$ for some $\kappa\in\mathbb{R}$. Then, $(x,y)\rightrightarrows D_{\ast}F(x,y)$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$.
Proof
We have there is a neighborhood $U\times V$ of $(\bar{x},\bar{y}-\kappa\bar{x})$ such that (1) is satisfied on $\operatorname{gph}(F-\kappa\text{Id})\cap(U\times V)$. For $(x,y)\in\operatorname{gph}F\cap\mathsf{T}_{\kappa}(U\times V)$, taking any $d\in D_{\ast}F(x,y)(z)$ for $z\in\mathcal{H}$, then there are sequences $(s_{k})\subset\mathbb{R}_{+}$, $(z_{k},d_{k})$ and $(x_{k},y_{k})\in\operatorname{gph}F$ such that $s_{k}\searrow 0$, $(z_{k},d_{k})\to(z,d)$, $(x_{k},y_{k})\to(x,y)$ and $(x_{k},y_{k})+s_{k}(z_{k},d_{k})\in\operatorname{gph}F$. Then, for all $k\in\mathbb{N}$ big enough, we have
$\langle(y_{k}+s_{k}d_{k})-y_{k},(x_{k}+s_{k}z_{k})-x_{k}\rangle\geq\kappa\|x_{k}+s_{k}z_{k}-x_{k}\|^{2}.$
It follows that $\langle d_{k},z_{k}\rangle\geq\kappa\|z_{k}\|^{2}$, then taking the limit when $k\to\infty$, we get $\langle d,z\rangle\geq\kappa\|z\|^{2}$, so we conclude that $D_{\ast}F(x,y)$ is $\kappa$-lower-definite.
An important example of locally maximal hypomonotone operator is given in the following proposition.
Proposition 3
Consider a set-valued mapping $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ with $(\bar{x},\bar{y})\in\operatorname{gph}F$. Suppose that $F^{-1}$ is strongly metrically regular at $(\bar{y},\bar{x})$. Then, $F$ is locally maximal hypomonotone.
Proof
Since $F^{-1}$ is strongly metrically regular at $(\bar{y},\bar{x})$, we have there is a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ and a Lipschitz function $\theta\colon U\to V$ such that $\operatorname{gph}F\cap U\times V=\operatorname{gph}\theta$. Let $\sigma>0$ be the Lipschitz constant of $\theta$, then for all $x,y\in U$,
$\langle\theta(y)-\theta(x),x-y\rangle\leq\|\theta(x)-\theta(y)\|\cdot\|x-y\|\leq\sigma\|x-y\|^{2}.$
It follows directly that $F+\sigma\text{Id}$ is locally monotone at $(\bar{x},\bar{y}+\sigma\bar{x})$. Regarding (3), consider $W:=\mathsf{T}_{\sigma}(U\times V)$, which is an open neighborhood of $(\bar{x},\bar{y}+\sigma\bar{x})$. Let us take a monotone operator $S\colon\mathcal{H}\rightrightarrows\mathcal{H}$ such that $W\cap\operatorname{gph}(F+\sigma\text{Id})\subset\operatorname{gph}S$. Then, consider $(x,y)\in W\cap\operatorname{gph}S$, we have $(x,y-\sigma x)\in U\times V$. Since $x\in U$, there is $\delta>0$ such that $\mathbb{B}_{\delta}(x)\subset U$, then for all $v\in\mathbb{B}$ and $\varepsilon\in[0,\delta[$, $x+\varepsilon v\in U$. Note that for all $z\in U$, $(z,\theta(z)+\sigma z)\in W\cap\operatorname{gph}S$, and given that $S$ is monotone, for all $v\in\mathbb{B}$ and $\varepsilon\in]0,\delta]$
$\langle x+\varepsilon v-x,\theta(x+\varepsilon v)+\sigma(x+\varepsilon v)-y\rangle\geq 0.$
It follows that
$\langle v,\theta(x+\varepsilon v)+\sigma\varepsilon v-(y-\sigma x)\rangle\geq 0.$
Then, taking $\varepsilon\searrow 0$, by virtue of the continuity of $\theta$, we have
$\langle v,\theta(x)-(y-\sigma x)\rangle\geq 0,\ \forall v\in\mathbb{B},$
and it implies that $\theta(x)=y-\sigma x$, so $(x,y-\sigma x)\in\operatorname{gph}F$, thus $(x,y)\in\operatorname{gph}(F+\sigma\text{Id})$. It follows that $W\cap\operatorname{gph}S=W\cap\operatorname{gph}(F+\sigma\text{Id})$, then $F+\sigma\text{Id}$ is locally maximal monotone at $(\bar{x},\bar{y}+\sigma\bar{x})$, concluding the desired result.
Lemma 5
Let $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ be a set-valued mapping and $(\bar{x},\bar{y})\in\operatorname{gph}F$. Suppose that $F$ is locally maximal hypomonotone at $(\bar{x},\bar{y})$, then $\operatorname{gph}F$ is locally closed at $(\bar{x},\bar{y})$.
Proof
Let us consider $\kappa>0$, and a neighborhood $U\times V$ of $(\bar{x},\bar{y}+\kappa\bar{x})$ such that $B:=\operatorname{gph}(F+\kappa\text{Id})\cap U\times V$ is monotone, and we also consider a maximal monotone extension of $B$, denoted by $G$ (see (MR3616647, , Theorem 20.21)). The local maximality of $F$ says us
$\operatorname{gph}(F+\kappa\text{Id})\cap U\times V=\operatorname{gph}G\cap U\times V.$
Since $G$ is maximal monotone (globally), we have $\operatorname{gph}G$ is closed (see (MR3616647, , Proposition 20.38 (iii))). Taking $W:=\mathsf{T}_{\kappa}^{-1}(U\times V)$, which is a neighborhood of $(\bar{x},\bar{y})$, we have
$\operatorname{gph}F\cap W=\mathsf{T}_{\kappa}^{-1}(\operatorname{gph}G)\cap W.$
Since $\mathsf{T}_{\kappa}^{-1}(\operatorname{gph}G)$ is closed, it follows $\operatorname{gph}F$ is locally closed at $(\bar{x},\bar{y})$.
3 Characterization of Local Strong Maximal Monotonicity
The theorem presented below corresponds to a characterization of strongly local maximal monotonicity through first-order properties related with lower definiteness of graphical derivatives (see Definition 2). This result can be seen as a counterpart criteria through coderivatives (see, e.g., (MR3485980, , Theorem 3.4)). First, we begin with an important lemma, which will be useful later.
Lemma 6
Let $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ be a set-valued mapping such that $DF$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})\in\operatorname{gph}F$ with $\kappa>0$ and $F$ is strongly metrically regular at $(\bar{x},\bar{y})$. Then,
  1. (a)
    $F$ is strongly locally monotone at $(\bar{x},\bar{y})$ with modulus $\kappa$.
  2. (b)
    $\text{reg}(F;(\bar{x},\bar{y}))\leq\frac{1}{\kappa}$.
Proof
Since $F$ is strongly metrically regular, there are a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ and a Lipschitz function $\theta\colon V\to U$ where $\operatorname{gph}F^{-1}\cap V\times U=\operatorname{gph}\theta$. By shrinking $V$, assume it is open and convex. Also, suppose that $DF(x,y)$ is $\kappa$-lower-definite for all $(x,y)\in U\times V\cap\operatorname{gph}F$.
Claim 1: For every $y\in V$ and $(u,v)\in\operatorname{gph}D\theta$ we have $\langle v,u\rangle\geq\|v\|^{2}$.
Proof of Claim 1: Take $y\in V$. Observe that,
$DF(\theta(y),y)^{-1}=DF^{-1}(y,\theta(y))=D\theta(y).$
Since $DF$ is $\kappa$-lower-definite it follows $\langle u,v\rangle\geq\kappa\|v\|^{2}$.
Claim 2: $F$ is strongly locally monotone at $(\bar{x},\bar{y})$ with modulus $\kappa$.
Proof of Claim 2: Take $y_{1},y_{2}\in V$. Define $\alpha\colon t\mapsto ty_{1}+(1-t)y_{2}$, we have that $\alpha([0,1])\subset V$ and in view of the Lipschitz continuity of $\theta$, we have $\theta\circ\alpha$ is absolutely continuous. By Lemma 2, we also have $D\theta(\alpha(t))(y_{2}-y_{1})=\{\frac{d}{dt}(\theta\circ\alpha)(t)\}$ for a.e. $t\in[0,1]$, thus Claim 1 yields
$\left\langle\frac{d}{dt}(\theta\circ\alpha)(t),y_{2}-y_{1}\right\rangle\geq\kappa\left\|\frac{d}{dt}(\theta\circ\alpha)(t)\right\|^{2}\text{ for a.e. }t\in[0,1].$
Then,
$\kappa\|\theta(y_{2})-\theta(y_{1})\|^{2}$ $=\kappa\|(\theta\circ\alpha)(1)-(\theta\circ\alpha)(0)\|^{2}$
[…]
$=\langle(\theta\circ\alpha)(1)-(\theta\circ\alpha)(0),y_{2}-y_{1}\rangle$
$=\langle\theta(y_{2})-\theta(y_{1}),y_{2}-y_{1}\rangle,$
where Proposition 1 was used, joint with Hölder inequality. Now, consider $(x_{i},y_{i})\in\operatorname{gph}F\cap U\times V$ for $i\in\{1,2\}$. It follows that $x_{i}=\theta(y_{i})$ for $i\in\{1,2\}$, then by the recently proved inequality, we have $\langle x_{1}-x_{2},y_{1}-y_{2}\rangle\geq\kappa\|x_{1}-x_{2}\|^{2}$.
Claim 3: $\text{reg}(F;(\bar{x},\bar{y}))\leq\frac{1}{\kappa}$.
Proof of Claim 3: Take $y_{1},y_{2}\in V$, by the inequality proved in Claim 2 for $\theta$ and Cauchy-Schwarz inequality, we have directly $\theta$ is $\frac{1}{\kappa}$-Lipschitz on $V$, implying that $\text{reg}(F;(\bar{x},\bar{y}))\leq\frac{1}{\kappa}$.
Now, we present our main result. It provides diverse characterizations describing local maximal strong monotonicity through graphical derivatives and some additional hypotheses.
Theorem 3.1
Let $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ be a set-valued mapping, and $(\bar{x},\bar{y})\in\operatorname{gph}F$. Suppose that $\kappa>0$. The following assertions are equivalent
  1. (a)
    $F$ is strongly locally maximal monotone at $(\bar{x},\bar{y})$ with modulus $\kappa$.
  2. (b)
    $DF$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$ and $F$ is locally maximal hypomonotone at $(\bar{x},\bar{y})$.
  3. (c)
    $DF$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$ and $F$ is strongly metrically regular at $(\bar{x},\bar{y})$.
Moreover, when $\mathcal{H}$ is finite dimensional, all the assertions are equivalent to
  1. (d)
    $D_{\ast}F$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$, $\operatorname{gph}F$ is locally closed at $(\bar{x},\bar{y})$ and $\bar{x}\in\liminf_{y\to\bar{y}}F^{-1}(y)$.
Proof
The proof of the equivalence of (a), (b) and (c) will be completed by demonstrating (a)$\Rightarrow$(b), (b)$\Rightarrow$(c) and (c)$\Rightarrow$(a).
(a)$\Rightarrow$(b): The lower definiteness of $DF$ follows from Lemma 4 and the local maximal hypomonotonicity is direct.
(b)$\Rightarrow$(c): We have $F$ is locally maximal hypomonotone at $(\bar{x},\bar{y})$, it follows that there is $\ell>0$ such that $F+\ell\text{Id}$ is locally maximal monotone at $(\bar{x},\bar{y}+\ell\bar{x})$. Then, by taking $\gamma>\ell$ fixed, from Proposition 2 we have that $F+\gamma\text{Id}$ is strongly locally maximal monotone at $(\bar{x},\bar{y}+\gamma\bar{x})$ with modulus $\gamma-\ell$. From (MR4769808, , Lemma 6.8) we have $F+\gamma\text{Id}$ is strongly metrically regular and by using Lemma 1 we have $D(F+\gamma\text{Id})$ is $(\kappa+\gamma)$-lower-definite around $(\bar{x},\bar{y}+\gamma\bar{x})$. Then, Lemma 6 says us $\text{reg}(F+\gamma\text{Id};(\bar{x},\bar{y}+\gamma\bar{x}))\leq\frac{1}{\gamma+\kappa}$. Define $g:=-\gamma\text{Id}$, we have that $g$ is $\gamma$-Lipschitz. Since $\frac{\gamma}{\kappa+\gamma}<1$, from (MR3288139, , Lemma 5F.1) we obtain directly that $F=(F+\gamma\text{Id})+g$ is strongly metrically regular at $(\bar{x},\bar{y})$ and we have done.
(c)$\Rightarrow$(a): By Lemma 6, we have $F$ is strongly locally monotone at $(\bar{x},\bar{y})$ with modulus $\kappa$. Since $F$ is also strongly metrically regular, the maximality follows from (MR4769808, , Lemma 6.8).
Now, we show the equivalence of (d) with the other assertions provided $\mathcal{H}$ is finite-dimensional.
(a)$\Rightarrow$(d): From Lemma 4, we have directly that $D_{\ast}F$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$. We use Lemma 5 to get that $\operatorname{gph}F$ is locally closed at $(\bar{x},\bar{y})$ and finally, the condition $\bar{x}\in\liminf_{y\to\bar{y}}F^{-1}(y)$ can be obtained directly since $F$ is strongly metrically regular (see (MR4769808, , Lemma 6.8)), then the assertion (d) is verified.
(d)$\Rightarrow$(b): Since $\operatorname{gph}DF(x,y)\subset\operatorname{gph}D_{\ast}F(x,y)$ for all $(x,y)\in\operatorname{gph}F$, we have directly that $DF$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$. We are going to prove that $F$ is strongly metrically regular at $(\bar{x},\bar{y})$. For one side, note that if $b\in\mathbb{B}$ and $a\in D_{\ast}F(\bar{x},\bar{y})^{-1}(b)$ we have $b\in D_{\ast}F(\bar{x},\bar{y})(a)$, then the lower-definiteness implies that $\kappa\|a\|^{2}\leq\langle a,b\rangle$. The Cauchy-Schwarz inequality yields $\|a\|\leq\frac{1}{\kappa}$, so it proves that
$|D_{\ast}F(\bar{x},\bar{y})^{-1}|^{+}:=\sup_{b\in\mathbb{B}}\sup_{a\in D_{\ast}F(\bar{x},\bar{y})^{-1}(b)}\|a\|<\infty,$
and together with the local closedness of $\operatorname{gph}F$ and $\bar{x}\in\liminf_{y\to\bar{y}}F^{-1}(y)$, we have $F$ is strongly metrically regular at $(\bar{x},\bar{y})$ (see (MR3288139, , Theorem 4D.1) which is especially true in finite-dimensional setting), then it implies that (b) holds and we have completed the proof.
Remark 1
Regarding the equivalence between (a) and (b) in the theorem above, as we previously pointed out, an analogous result based on coderivatives is given in (MR3485980, , Theorem 3.4). In contrast to our result, note that we require local maximal hypomonotonicity in (b), whereas their result does not impose maximality. In Example 1 below, we demonstrate that this assumption is necessary in our case.
The following example shows that the statements (b), (c) and (d) in Theorem 3.1 cannot be weakened in order to only require the lower definiteness of $DF$ (or $D_{\ast}F$) and to obtain the same conclusion.
Example 1
Consider $F\colon\mathbb{R}\rightrightarrows\mathbb{R}$ given by $F(x)=\{x\}$ for $x<0$, $F(x)=\{x+1\}$ for $x>0$ and $F(0)=\{0,1\}$. Note that for all $(x,y)\in\operatorname{gph}F$ and $(u,v)\in\operatorname{gph}D_{\ast}F(x,y)$, $\langle u,v\rangle\geq\|u\|^{2}$. However, note that $0\notin\liminf_{y\to 0}F^{-1}(y)$. We can also see that $F$ is not strongly metrically regular at $(0,0)$. Finally, we can see that $F$ is strongly locally monotone at $(0,0)$ but the maximality fails.
In the above example the property (c) fails due to $F$ is not strongly metrically regular, but it is not metrically regular either. Then, it is worth asking if it is possible to change strong metric regularity in (c) by a weaker condition. Observe that strong metric subregularity (in finite dimension) is not enough since Lemma 3. The following example shows that the assumption of metric regularity is not enough either.
Example 2
Consider $F\colon\mathbb{R}\rightrightarrows\mathbb{R}$ given by $F(x)=\{x\}$ for $x<0$ and $F(x)=\{x,2x\}$ for $x\geq 0$. At $(0,0)$, $F$ is not locally monotone and we can see that $DF(x,y)$ is $1$-lower-definite for all $(x,y)\in\operatorname{gph}F$. We can also prove that $\forall x,y\in\mathbb{R}$, $d(x;F^{-1}(y))\leq d(y;F(x))$, in particular, $F$ is metrically regular at $(0,0)$, but it is not strongly metrically regular there.
The following simple example shows that some equivalences stated in Theorem 3.1 are not satisfied when $\kappa\leq 0$.
Example 3
Take $F\colon\mathbb{R}\rightrightarrows\mathbb{R}$ given by $F(x)=\{0\}$ for all $x\in\mathbb{R}$. Note that $F$ is maximal monotone (globally) and $DF$ is 00-lower-definite at every point of $\operatorname{gph}F$. However, for all $(x_{0},y_{0})\in\operatorname{gph}F$, $x_{0}\notin\liminf_{y\to y_{0}}F^{-1}(y)$ and $F$ is not strongly metrically regular at any point of $\operatorname{gph}F$. It follows that (a)$\Rightarrow$(c) and (a)$\Rightarrow$(d) fail.
However, we can establish the equivalence of the assertions (a) and (b) in Theorem 3.1.
Corollary 1
Let $F\colon\mathcal{H}\rightrightarrows\mathcal{H}$ be a set-valued mapping, and $(\bar{x},\bar{y})\in F$. Let $\kappa\in\mathbb{R}$. Then $F-\kappa\text{Id}$ is locally maximal monotone at $(\bar{x},\bar{y}-\kappa\bar{x})$ if and only if $DF$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$ and $F$ is locally maximal hypomonotone at $(\bar{x},\bar{y})$.
Proof
The necessity part is straightforward from Lemma 4. So, now we are going to deal with the converse. Observe that the case $\kappa>0$ is done in Theorem 3.1, so we suppose that $\kappa\leq 0$. Since $F$ is locally maximal hypomonotone at $(\bar{x},\bar{y})$, there is $\ell>0$ such that $F+\ell\text{Id}$ is locally maximal monotone at $(\bar{x},\bar{y}+\ell\bar{x})$. From Proposition 2, we can suppose that $\ell>-\kappa$. From Lemma 1 we have $F+\ell\text{Id}$ is $(\ell+\kappa)$-lower-definite around $(\bar{x},\bar{y}+\ell\bar{x})$, so we can apply Theorem 3.1 in order to obtain that $F+\ell\text{Id}$ is strongly locally maximal monotone at $(\bar{x},\bar{y}+\ell\bar{x})$ with modulus $\ell+\kappa$. By using Proposition 2, we have $F-\kappa\text{Id}$ is locally maximal monotone at $(\bar{x},\bar{y}-\kappa\bar{x})$, concluding the desired.
4 Characterization of Variational Convexity
This section is devoted to study how acts Theorem 3.1 when the set-valued mapping is the subdifferential of a given function. We are going to prove that it allows to give a second-order characterization of variational convexity, based on first-order analysis of subdifferential mapping. This kind of results has been recently studied in gfrerer2025 in finite-dimensional setting, by performing different techniques to the ones presented in our work. Furthermore, in MR4784083 are presented characterizations of variational convexity through first order properties of coderivatives, where is assumed subdifferential continuity (see Definition 4). First, we recall some definitions and properties from variational analysis.
For a given set $S\subset\mathcal{H}$, we define its indicator function as $x\mapsto\delta(x;S)$ where $\delta(x;S)=0$ if $x\in S$ and $\delta(x;S)=\infty$ if $x\notin S$.
Consider a function $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$ and $x\in\operatorname{dom}f$. We say an element $\zeta$ belongs to the proximal subdifferential of $f$ at $x$, denoted by $\partial_{P}f(x)$ if there are $\eta,\delta>0$ such that
$f(y)\geq f(x)+\langle\zeta,y-x\rangle-\frac{\sigma}{2}\|y-x\|^{2},\forall y\in\mathbb{B}_{\delta}(x).$
Moreover, an element $\zeta$ belongs to the limiting subdifferential of $f$ at $x$, denoted by $\partial f(x)$, if there are sequences $(x_{n})\to x$ and $(\zeta_{n})\rightharpoonup\zeta$ such that $f(x_{n})\to f(x)$ and $\forall n\in\mathbb{N}:\zeta_{n}\in\partial_{P}f(x_{n})$.
A given function $f\colon\mathcal{H}\to\overline{\mathbb{R}}$ is said to be $\sigma$-convex for some $\sigma\in\mathbb{R}$ if $f-\frac{\sigma}{2}\|\cdot\|^{2}$ is convex on $\mathcal{H}$. We also say $f$ is locally $\sigma$-convex at $\bar{x}\in\operatorname{dom}f$ if there is a convex open neighborhood $U$ of $\bar{x}$ such that $f+\delta(\cdot;U)$ is $\sigma$-convex. When $\sigma=0$, we simply omit the dependence of $\sigma$.
Definition 3
Let $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$ be a proper function. Given $\bar{x}\in\operatorname{dom}\partial f$ and $\bar{y}\in\partial f(\bar{x})$, we say $f$ is prox-regular at $\bar{x}$ for $\bar{y}$ if there are some reals $\varepsilon,\sigma>0$ such that for all $x,x^{\prime}\in\mathbb{B}_{\varepsilon}(\bar{x})$ with $|f(x)-f(\bar{x})|<\varepsilon$ and all $y\in\mathbb{B}_{\varepsilon}(\bar{y})$ with $y\in\partial f(x)$ we have
$f(x^{\prime})\geq f(x)+\langle y,x^{\prime}-x\rangle-\frac{\sigma}{2}\|x^{\prime}-x\|^{2}.$ (4)
Definition 4
A proper function $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$ is said to be subdifferentially continuous at $\bar{x}$ for $\bar{y}\in\partial f(\bar{x})$ if $(x,v)\mapsto f(x)$ is continuous at $(\bar{x},\bar{y})$ relative to $\operatorname{gph}\partial f$.
Definition 5
An lsc function $f$ will be called variationally convex at $\bar{x}$ for $\bar{y}\in\partial f(\bar{x})$ with modulus $\sigma\in\mathbb{R}$ if for some open convex neighborhood $U\times V$ of $(\bar{x},\bar{y})$ there is a $\sigma$-convex lsc function $\hat{f}\leq f$ on $U$ such that, for some $\varepsilon>0$,
$U_{\varepsilon}\times V\cap\operatorname{gph}\partial f=U\times V\cap\operatorname{gph}\partial\hat{f}$
and
$\quad f(x)=\hat{f}(x)\text{ for all }(x,y)\in U_{\varepsilon}\times V\cap\operatorname{gph}\partial f.$
where $U_{\varepsilon}:=\{x\in U:f(x)<f(\bar{x})+\varepsilon\}$. When $\sigma=0$ we simply say $f$ is variationally convex at $\bar{x}$ for $\bar{y}$.
Lemma 7
Consider a proper and lsc function $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$. Given $\bar{x}\in\operatorname{dom}f$, $\bar{y}\in\partial f(\bar{x})$ and $\sigma\in\mathbb{R}$, one has the following assertions are equivalent
  1. (a)
    $f$ is variationally convex at $\bar{x}$ for $\bar{y}$ with modulus $\sigma$.
  2. (b)
    $f_{\sigma}:=f-\frac{\sigma}{2}\|\cdot-\bar{x}\|^{2}$ is variationally convex at $\bar{x}$ for $\bar{y}$.
Proof
Suppose (a) holds. When $\sigma=0$ there is nothing to prove. Suppose then, $\sigma\neq 0$. Then, there is an open convex neighborhood $U\times V$ of $(\bar{x},\bar{y})$ and a $\sigma$-convex function $\hat{f}\leq f$ on $U$ such that, for some $\varepsilon>0$,
$U_{\varepsilon}^{f}\times V\cap\operatorname{gph}\partial f=U\times V\cap\operatorname{gph}\partial\hat{f},$ (5)
with $f=\hat{f}$ on the points where (5) is satisfied. Define $\hat{f}_{\sigma}:=\hat{f}-\frac{\sigma}{2}\|\cdot-\bar{x}\|^{2}$, we have $\hat{f}_{\sigma}$ is convex. The sum rule for limiting subdifferential says us $\partial f_{\sigma}(x)=\partial f(x)-\sigma(x-\bar{x})$ for all $x\in\operatorname{dom}f$. Define $\Pi_{\sigma}\colon x\mapsto(x,y+\sigma(x-\bar{x}))$, then, there is $\delta>0$ such that $W\times Z\subset\Pi_{\sigma}^{-1}(U\times V)$ where $W=\mathbb{B}_{\delta}(\bar{x})$ and $Z=\mathbb{B}_{\delta}(\bar{y})$, then by (5) we have
$W_{\varepsilon}^{f}\times Z\cap\operatorname{gph}\partial f_{\sigma}=W\times Z\cap\operatorname{gph}\partial\hat{f}_{\sigma}.$
Consider $\mathcal{W}=\mathbb{B}_{r}(\bar{x})$ where $r=\min\{\delta,\sqrt{\varepsilon/|\sigma|}\}$. The choice of $r$ implies
$\mathcal{W}_{\varepsilon/2}^{f_{\sigma}}=\{x\in\mathcal{W}:f_{\sigma}(x)<f_{\sigma}(\bar{x})+\varepsilon/2\}\subset\{x\in\mathcal{W}:f(x)<f(\bar{x})+\varepsilon\}.$
Finally, we can verify that $\mathcal{W}_{\varepsilon/2}^{f_{\sigma}}\times Z\cap\operatorname{gph}f_{\sigma}=\mathcal{W}\times Z\cap\operatorname{gph}\partial\hat{f}_{\sigma}$ and it is easy to see that $f_{\sigma}=\hat{f}_{\sigma}$ there. Then, (b) holds. The converse is quite similar to the argument just given.
For $\varepsilon>0$, we denote by $\mathscr{F}_{\varepsilon}^{f}$ to the $f$-attentive localization of $\partial f$ at $(x_{0},y_{0})\in\operatorname{gph}\partial f$ given by
$\mathscr{F}_{\varepsilon}^{f}\colon x\rightrightarrows\{y:y\in\partial f(x),\|x-x_{0}\|<\varepsilon,\|y-y_{0}\|<\varepsilon,f(x)<f(x_{0})+\varepsilon\}.$
Proposition 4
Consider a proper and lsc function $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$. Let $x_{0}\in\operatorname{dom}f$ and $y_{0}\in\partial f(x_{0})$. Then, the following assertions are equivalent
  1. (a)
    $f$ is prox regular at $x_{0}$ for $y_{0}$.
  2. (b)
    $y_{0}\in\partial_{P}f(x_{0})$ and there is $\bar{\varepsilon}>0$, such that for all $\varepsilon\in]0,\bar{\varepsilon}[$, $\mathscr{F}_{\varepsilon}^{f}$ is locally maximal hypomonotone at $(x_{0},y_{0})$.
  3. (c)
    $y_{0}\in\partial_{P}f(x_{0})$ and $\mathscr{F}_{\varepsilon}^{f}$ is locally hypomonotone at $(x_{0},y_{0})$ for some $\varepsilon>0$.
Proof
First, note that (b)$\Rightarrow$(c) is straightforward and (c)$\Rightarrow$(a) follows from (MR4659163, , Theorem 11.50). It remains to prove (a)$\Rightarrow$(b). Indeed, suppose $f$ is prox regular at $x_{0}$ for $y_{0}$ with parameters $\hat{\varepsilon}>0$ and $\hat{\sigma}>0$. From (4) follows directly that $y_{0}\in\partial_{P}f(x_{0})$. By using (4) again, it is clear that for all $\varepsilon\in]0,\hat{\varepsilon}[$, $\mathscr{F}_{\varepsilon}^{f}$ is locally hypomonotone at $(x_{0},y_{0})$. Since $f$ is lsc at $x_{0}$, there is $\eta>0$ such that $f$ is bounded from below on $\mathbb{B}_{\eta}[x_{0}]$, then defining $\bar{f}=f+\delta(\cdot;\mathbb{B}_{\eta}[x_{0}])$ we have $\operatorname{gph}\mathscr{F}_{\varepsilon}^{f}=\operatorname{gph}\mathscr{F}_{\varepsilon}^{\bar{f}}$ for all $\varepsilon<\min\{\eta,\hat{\varepsilon}\}=:\bar{\varepsilon}$. Note that $\bar{f}$ is prox regular at $x_{0}$ for $y_{0}$ and it is also prox bounded. Consider any $\varepsilon<\bar{\varepsilon}$, then by virtue of (MR4659163, , Lemma 11.51), for all $\lambda>0$ small enough, $(\text{Id}+\lambda\mathscr{F}_{\varepsilon}^{\bar{f}})^{-1}$ is single-valued and continuous on a neighborhood of $x_{0}+\lambda y_{0}$. From (MR4769808, , Theorem 8.5) it follows that $\mathscr{F}_{\varepsilon}^{\bar{f}}$ is locally maximal hypomonotone at $(x_{0},y_{0})$, concluding the desired implication.
Given $\bar{x}\in\mathcal{H}$, for a neighborhood $U$ of $\bar{x}$, $\varepsilon>0$ and $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$, we define
$U_{\varepsilon}^{f}=\{x\in U:f(x)<f(\bar{x})+\varepsilon\}.$
We also define the $f$-attentive graphical derivative (introduced in gfrerer2025 ) of $f$ at $(x,y)\in\operatorname{gph}\partial f$ in the direction $u$ as
$D_{f}(\partial f)(x,y)(u)=\left\{v\;\middle|\;\begin{aligned} &\exists s_{k}\searrow 0,\exists(u_{k},v_{k})\to(u,v)\text{ s.t. }\forall k\in\mathbb{N},\\ &y+s_{k}v_{k}\in\partial f(x+s_{k}u_{k})\text{ and }f(x+s_{k}u_{k})\to f(x)\end{aligned}\right\}.$
Observe that when $f$ is subdifferentially continuous, $D_{f}(\partial f)$ coincides with $D(\partial f)$.
We recall that a function $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$ is said to be sequentially weakly lsc at $x\in\operatorname{dom}f$ if for all $(x_{n})\rightharpoonup x$, $\liminf_{n\to\infty}f(x_{n})\geq f(x)$.
The following theorem can be seen as the infinite-dimensional counterpart of the equivalence (i)$\iff$(iv) given in (gfrerer2025, , Theorem 4.3). Our approach consists in applying Theorem 3.1 and Corollary 1, regarding Proposition 4 since locally maximal hypomonotonicity is required.
Theorem 4.1
Let us consider a proper function $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$ with $\bar{x}\in\operatorname{dom}f$, $\bar{y}\in\partial f(\bar{x})$ and $f$ is sequentially weakly lsc around $\bar{x}$. Let us consider $\kappa\in\mathbb{R}$. Then, the following statements are equivalent
  1. (a)
    $f$ is variationally convex at $\bar{x}$ for $\bar{y}$ with modulus $\kappa$.
  2. (b)
    $f$ is prox regular at $\bar{x}$ for $\bar{y}$ and there exist $\varepsilon>0$ and a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ such that $D_{f}(\partial f)$ is $\kappa$-lower-definite on $U_{\varepsilon}^{f}\times V$.
Proof
(a)$\Rightarrow$(b): The prox regularity of $f$ is immediate since $f$ is variationally convex. For other side, there exist $\varepsilon>0$, a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ and a $\kappa$-convex and lsc function $\varphi\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$ such that $f\geq\varphi$ on $U$,
$U_{\varepsilon}^{f}\times V\cap\operatorname{gph}\partial f=U\times V\cap\operatorname{gph}\partial\varphi\text{ and }f=\varphi\text{ on }U_{\varepsilon}^{f}\times V\cap\operatorname{gph}\partial f.$
Note that if $(x,y)\in U_{\varepsilon}^{f}\times V\cap\operatorname{gph}\partial f$ and $(u,v)\in\operatorname{gph}D_{f}(\partial f)(x,y)$, there are $(u_{k},v_{k})\to(u,v)$ and $s_{k}\searrow 0$ such that $y+s_{k}v_{k}\in\partial f(x+s_{k}u_{k})$ for all $k\in\mathbb{N}$ and $f(x+s_{k}u_{k})\to f(x)$. Then, for all $k\in\mathbb{N}$ big enough,
$(x+s_{k}u_{k},y+s_{k}v_{k})\in U_{\varepsilon}^{f}\times V\cap\operatorname{gph}\partial f.$
The $\kappa$-convexity of $\varphi$ implies that for all $k\in\mathbb{N}$ big enough
$\langle x+s_{k}u_{k}-x,y+s_{k}v_{k}-y\rangle\geq\kappa\|x+s_{k}u_{k}-x\|^{2},$
it follows that $\langle u_{k},v_{k}\rangle\geq\kappa\|u_{k}\|^{2}$, then taking $k\to\infty$ we conclude that $\langle u,v\rangle\geq\kappa\|u\|^{2}$ and it allows to conclude that $D_{f}(\partial f)$ is $\kappa$-lower-definite on $U_{f}^{\varepsilon}\times V$.
(b)$\Rightarrow$(a): Since $f$ is prox regular, we have there is $\bar{\varepsilon}>0$ such that for all $\eta\in]0,\bar{\varepsilon}[$ the following localization
$\mathscr{F}_{\eta}^{f}\colon x\rightrightarrows\{y:y\in\partial f(x),\|x-\bar{x}\|<\eta,\|y-\bar{y}\|<\eta,f(x)<f(\bar{x})+\eta\}$
is locally maximal hypomonotone at $(\bar{x},\bar{y})$ (see Proposition 4). Moreover, suppose (4) holds for some $\varepsilon^{\prime}>0$ and $\sigma>0$. From now on, we fix $\eta<\min\{\bar{\varepsilon},\varepsilon,\varepsilon^{\prime}\}$ such that $\mathbb{B}_{\eta}(\bar{x})\times\mathbb{B}_{\eta}(\bar{y})\subset U\times V$. We split this part by proving the following claims:
Claim 1: For all $(x,y)\in\mathbb{B}_{\eta}(\bar{x})\times\mathbb{B}_{\eta}(\bar{y})$, $\operatorname{gph}D\mathscr{F}_{\eta}^{f}(x,y)\subset\operatorname{gph}D_{f}(\partial f)(x,y)$.
Proof of Claim 1: Indeed, let $(u,v)\in\operatorname{gph}D\mathscr{F}_{\eta}^{f}(x,y)$, then there are sequences $s_{k}\searrow 0$ and $(u_{k},v_{k})\to(u,v)$ such that for all $k\in\mathbb{N}$, $y+s_{k}v_{k}\in\mathscr{F}_{\eta}^{f}(x+s_{k}u_{k})$. Note that it implies $y+s_{k}v_{k}\in\partial f(x+s_{k}u_{k})$, $f(x+s_{k}u_{k})<f(\bar{x})+\eta$, $x+s_{k}u_{k}\in\mathbb{B}_{\eta}(\bar{x})$ and $y+s_{k}v_{k}\in\mathbb{B}_{\eta}(\bar{y})$ for all $k\in\mathbb{N}$. Then, regarding (4), we have for all $k\in\mathbb{N}$
$f(x)\geq f(x+s_{k}u_{k})+\langle y+s_{k}v_{k},x-(x+s_{k}u_{k})\rangle-\frac{\sigma}{2}\|x-(x+s_{k}u_{k})\|^{2}.$
It follows that,
$f(x)\geq f(x+s_{k}u_{k})-s_{k}\langle y+s_{k}v_{k},u_{k}\rangle-\frac{\sigma s_{k}^{2}}{2}\|u_{k}\|^{2}.$
Taking $k\to\infty$ we conclude that $\limsup_{k\to\infty}f(x+s_{k}u_{k})\leq f(x)$ and since $f$ is lsc we conclude $f(x)=\lim_{k\to\infty}f(x+s_{k}u_{k})$, it follows that $(u,v)\in\operatorname{gph}D_{f}(\partial f)(x,y)$.
Claim 2: $D\mathscr{F}_{\eta}^{f}$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$.
Proof of Claim 2: Define $\mathcal{U}:=\mathbb{B}_{\eta}(\bar{x})$ and $\mathcal{V}:=\mathbb{B}_{\eta}(\bar{y})$. Note that, Claim 1 implies that $D\mathscr{F}_{\eta}^{f}$ is $\kappa$-lower-definite on $\mathcal{U}_{\varepsilon}^{f}\times\mathcal{V}$. Since $\eta<\varepsilon$, we have $\operatorname{dom}\mathscr{F}_{\eta}^{f}\subset\mathcal{U}_{\varepsilon}^{f}$. It follows that $D\mathscr{F}_{\eta}^{f}$ is $\kappa$-lower-definite on $\mathcal{U}\times\mathcal{V}$.
Claim 3: For $\kappa\geq 0$, (b)$\Rightarrow$(a) holds.
Proof of Claim 3: Suppose $\kappa>0$. Since $\mathscr{F}_{\eta}^{f}$ is locally maximal hypomonotone at $(\bar{x},\bar{y})$ and $D\mathscr{F}_{\eta}^{f}$ is $\kappa$-lower-definite around $(\bar{x},\bar{y})$, by using Theorem 3.1 we have directly that $\mathscr{F}_{\eta}^{f}$ is strongly locally maximal monotone with modulus $\kappa$. From (MR4784083, , Theorem 6.10), since $f$ is also weakly sequentially lsc around $\bar{x}$, we conclude $f$ is variationally convex at $\bar{x}$ for $\bar{y}$ with modulus $\kappa$. When $\kappa=0$, we proceed similarly, concluding by Corollary 1 that $\mathscr{F}_{\eta}^{f}$ is locally maximal monotone at $(\bar{x},\bar{y})$. Finally, by (MR4784083, , Theorem 6.6), we conclude $f$ is variationally convex at $\bar{x}$ for $\bar{y}$.
Claim 4: (b)$\Rightarrow$(a) holds for $\kappa<0$.
Proof of Claim 4: Consider $g:=f-\frac{\kappa}{2}\|\cdot-\bar{x}\|^{2}$. First of all, note that since $f$ is weakly sequentially lsc around $\bar{x}$ and $\frac{-\kappa}{2}\|\cdot-\bar{x}\|^{2}$ is convex and continuous, the function $g$ is sequentially weakly lsc around $\bar{x}$. On the one hand, since $f$ is prox regular at $\bar{x}$ for $\bar{y}$, by virtue of Proposition 4 it is not difficult to see that $g$ is prox regular at $\bar{x}$ for $\bar{y}$. On the other hand, by the sum rule for limiting subdifferential, we have $\partial g(x)=\partial f(x)-\kappa(x-\bar{x})$ for all $x\in\operatorname{dom}f$. We consider $\Pi_{\kappa}\colon(x,y)\mapsto(x,y+\kappa(x-\bar{x}))$. Consider $\delta>0$ such that $\mathbb{B}_{\delta}(\bar{x})\times\mathbb{B}_{\delta}(\bar{y})\subset\Pi_{\kappa}^{-1}(U\times V)$ and define $\mathscr{U}:=\mathbb{B}_{r}(\bar{x})$ and $\mathscr{V}=\mathbb{B}_{\delta}(\bar{y})$. Consider $(x,y)\in\mathscr{U}_{\varepsilon}^{g}\times\mathscr{V}$ and $(u,v)\in\operatorname{gph}D_{g}(\partial g)(x,y)$, then
$(u,v+\kappa u)\in\operatorname{gph}D_{f}(\partial f)(x,y+\kappa(x-\bar{x}))\text{ and }(x,y+\kappa(x-\bar{x}))\in U\times V.$
We also have $x\in U_{\varepsilon}^{f}$ due to $\kappa<0$. Since $D_{f}(\partial f)$ is $\kappa$-lower-definite on $U_{\varepsilon}^{f}\times V$, we have $\langle u,v+\kappa u\rangle\geq\kappa\|u\|^{2}\Rightarrow\langle u,v\rangle\geq 0$. It follows that $D_{g}(\partial g)$ is 00-lower-definite on $\mathscr{U}_{\varepsilon}^{g}\times\mathscr{V}$. To finalize, note that $g$ satisfies the assumptions required to apply Claim 3 (with $\kappa=0$), then we have $g$ is variationally convex at $\bar{x}$ for $\bar{y}$, thus, applying Lemma 7, we conclude $f$ is variationally convex at $\bar{x}$ for $\bar{y}$ with modulus $\kappa$.
From the last result, we can derive a characterization of local convexity of a $\mathcal{C}^{1,+}$ function, i.e. a continuously differentiable function whose gradient is locally Lipschitz.
Corollary 2
Let us consider a function $f\colon\mathcal{H}\to\mathbb{R}\cup\{\infty\}$ which is of class $\mathcal{C}^{1,+}$ around a point $\bar{x}\in\operatorname{dom}f$ and also sequentially weakly lsc there. Let $\kappa\in\mathbb{R}$. Then, the following assertions are equivalent
  1. (a)
    $f$ is locally $\kappa$-convex at $\bar{x}$.
  2. (b)
    $D(\nabla f)$ is $\kappa$-lower-definite around $(\bar{x},\nabla f(\bar{x}))$.
Proof
The function $f$ is prox regular at $\bar{x}$ for $\nabla f(\bar{x})$, as a consequence of combining Proposition 3 and Proposition 4. Note that $f$ is also subdifferentially continuous at $\bar{x}$ for $\nabla f(\bar{x})$. Moreover, if $f$ is variationally convex at $\bar{x}$ for $\nabla f(\bar{x})$ with modulus $\kappa$, then the function $\bar{f}=f-\frac{\kappa}{2}\|\cdot-\bar{x}\|^{2}$ is variationally convex at $\bar{x}$ for $\nabla f(\bar{x})$ (Lemma 4.1), and it is still of class $\mathcal{C}^{1,+}$ around $\bar{x}$, then by (MR4769808, , Remark 8.15 (iv)) we have $\bar{f}$ is locally convex around $\bar{x}$ and therefore $f$ is locally $\kappa$-convex around $\bar{x}$. Then, Theorem 4.1 allows us to conclude the equivalence between (a) and (b).
5 Concluding Remarks
In this paper, we have established new characterizations of local maximal monotonicity for set-valued mappings based on the lower-definiteness of the graphical derivative. In particular, Theorem 3.1 provides three distinct characterizations of strong maximal monotonicity. Finally, in Theorem 4.1, we apply our main results to characterize variational convexity, leveraging its connection with the local maximal monotonicity properties of the subdifferential mapping. This can be seen as a second-order characterization of variational convexity, complementing the result in (MR4784083, , Theorem 7.3), where coderivatives are used and subdifferential continuity is required.
The results obtained strengthen the theory of locally monotone operators and offer potential applications in second-order theory, stability of solution mappings, and related areas.